Question:
Directions: Write a system of equations to model the situation.
Sarah is 25 years older than her son Gavin. In ten years, Sarah will be twice Gavin's age. How old are Sarah and Gavin now?
My Work:
Let x= Sarah's age
Let y= Gavin's age
y=x-25
10+x=2y
Subsitution:
y=2y-25
+25
y+25=2y
-y
25=2y
divide by 2
y=25
2(25)=x
x=50 Sarah is 50 years old and Gavin is 25 years old.

Answer by Alan3354(69443)

Question:
Directions: Write a system of equations to model the situation.
Sarah is 25 years older than her son Gavin. In ten years, Sarah will be twice Gavin's age. How old are Sarah and Gavin now?
My Work:
Let x= Sarah's age
Let y= Gavin's age
y=x-25
10+x=2y ************** In 10 years, 2*(x-15) = y+10
Subsitution:
y=2y-25
+25
y+25=2y
-y
25=2y
divide by 2
y=25
2(25)=x
x=50 Sarah is 50 years old and Gavin is 25 years old.
------------------------
The answer you got is what their ages will be in 10 years.
Now, they're 40 and 15.
=============================